A supermarket is selling 2 types of candies, orange slices and strawberry leaves. The orange slices cost $1.27 per pound and the strawberry leaves cost $1.77 per pound. How many pounds of each should be mixed to get a 9-pound mixture that sells for $12.93?

Answer by Boreal(15235)

x pounds of orange slices
9-x pounds of strawberry leaves
x(1.27)+(9-x)(1.77)=12.93
1.27x+15.93-1.77x=12.93
-0.5x=-3
x=6 pounds of orange slices ($7.62)
9-x=3 pounds of strawberry leaves ($5.31)
add to $12.93